A recent survey showed that among 100 randomly selected college seniors, 20 plan to attend graduate school and 80 do not. Determine a 68 % confidence interval for the population proportion of college seniors who plan to attend graduate school. (Enter each answer rounded to three decimal places.)
Answer by stanbon(75887)

A recent survey showed that among 100 randomly selected college seniors, 20 plan to attend graduate school and 80 do not. Determine a 68 % confidence interval for the population proportion of college seniors who plan to attend graduate school. (Enter each answer rounded to three decimal places.)
-----
p-hat = 20/100 = 1/5
------
ME = z*s = 0.9945*sqrt[(1/5)(4/5)/100] = 0.0398
-------
68% CI: 0.2-0.0398 < p < 0.2+0.0398
=========================================
cheers,
Stan H.
==================